Hanson drove 200 kilometers at x km/hr and then decreased his speed 20 km/hr and drove 240 extra kilometers. Write a rational expression for his total traveling time. Evaluate that expression for x = 100 km/hr.
Answer by mananth(16949)

200 km at xkm/h
time = 200/x
..
decreased speed = x-20
distance traveled = 240 km
time = 240 /(x-20)
..
Total time T = 200/x + 240/(x-20)
For x =100
..
T = 200/100 + 240/(80)
T=2+3 = 5 hours